According to given data First we need to calculate the means of each sample Sample1 = (9.5+8.4+9.8+10.3)/4 = 9.5 Sample 2 = (10.3+9.9+10.2+10.5)/4 = 10.225 Sample 3 = (9.9+10.5+9.9+10.8)/4 = 10.275 Sample 4 = (10.3+10.4+10.8+10.6)/4=10.525 Sample 5 (9.6+10.9+10.8+10.1)/4=10.35 Sample 6 (10.2+8.9+9.7+10.5)/4=9.825

Find standard deviation


1
Expert's answer
2021-11-25T02:00:03-0500

Sample1 = (9.5+8.4+9.8+10.3)/4 = 9.5 Sample 2 = (10.3+9.9+10.2+10.5)/4 = 10.225 Sample 3 = (9.9+10.5+9.9+10.8)/4 = 10.275 Sample 4 = (10.3+10.4+10.8+10.6)/4=10.525 Sample 5 (9.6+10.9+10.8+10.1)/4=10.35 Sample 6 (10.2+8.9+9.7+10.5)/4=9.825